OpenStudy (anonymous):

Can someone please help me by telling the formula to calculate inflation via CPI

OpenStudy (anonymous):

CPI at time t = 100 CPI at time t+1 = x inflation = (CPI at time t+1-CPI at t)/CPI at time t. * 100 = x-100 If x is say 106. inflation in time period t to t+1 has been 6%. Simple enough. Harder when you add multiple years, and then you can only take out the average inflation. Using the formula for compound interest: A = P(1+r/100)^t r/100 = (A/P)^1/t - 1 Substitute A = CPI at time y; P is base CPI and t is time period. R is thus average inflation over the time period.